OpenStudy (anonymous):

What is the value of w? 2(2w-6)=6(w-6)+14

OpenStudy (anonymous):

@ZAO1234

OpenStudy (anonymous):

@Rubyjadeluv

OpenStudy (anonymous):

w=5

OpenStudy (anonymous):

Did you want a step by step process?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

Okay

OpenStudy (jhannybean):

2(2w) + (-6 * 2) = 6(w) + (-6 * 6) +14 Simplify that first.

OpenStudy (anonymous):

\[2(2w-6)=6(w-6)+14\]\[4w-12=6w-36+14\]\[4w-6w=-36+12+14\]\[-2w=-24+14\]\[-2w=-10\]\[w=5\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome!